OpenStudy (anonymous):

How do I simplify (16x^3)(3x^4y)(x^5y^15)?

OpenStudy (angela210793):

=48*x^12*y^16

OpenStudy (anonymous):

exponent laws x^a * x^b = x^a+b

OpenStudy (anonymous):

\[48 x ^{12} y ^{16}\]